Determine whether the given quadratic function has a minimum value or maximum value. Then find the coordinates of the minimum or maximum point.
f(x) = -x squared - 2x - 5

Answer by oscargut(686)

if f(x)=ax^2+bx+c then V=(-b/2a,f(-b/2a))
then if f(x)=-x^2-2x-5 V=(-1,f(-1))=(-1,-4)
and it has a maximum value because starts with a negative coefficient